### v1

# Effects of dark boson mediated feeble interaction between dark matter (DM) and quark matter on $f$-mode oscillation of DM admixed quark stars

Abstract
We investigate the behaviour of the prominent non-radial fundamental $f$-mode oscillations of dark matter (DM) admixed strange quark stars (DMSQSs), by adopting an equation of state (EoS) developed in Ref. Sen and Guha (2022), which considers the possible presence of feebly interacting DM in strange quark stars (SQSs) for the first time. Within the model, feeble interaction between fermionic DM $\chi$ and strange quark matter (SQM) is invoked via a vector new physics mediator $\xi$ with coupling strength $y_{\xi}$. The pure SQM is described by the vector MIT Bag model. By varying different EoS parameters, the structural properties viz. the mass, radius and tidal deformability ($\Lambda$) of the DMSQSs are studied with respect to various astrophysical constraints. We study in detail the $f$-mode spectra within the Cowling approximation by obtaining the frequencies as a function of mass, compactness and $\Lambda$ of the star. To the best of our knowledge, this study represents the first analysis of non-radial $f$-mode oscillations of DMSQSs. Our investigation indicates that the presence of DM and its interaction with SQM has great impact on the $f$-modes. We show that the $f$-mode frequencies are larger for DMSQSs, which are largely populated with massive DM fermions, compared to the SQSs. Further, we obtain a linear empirical relation between the $f$-modes and the average density of the star. We also find that the mass-scaled angular frequency varies universally with compactness for DMSQSs.
I Introduction
One of the most interesting facts about compact stars is that they present the opportunity to study the matter at extreme conditions of density (about $5-10$ times the normal nuclear density), compactness and gravity. Their composition at such high density is still experimentally inconclusive and presently we adopt various theoretical modeling of the constituent matter equation of state (EoS) to understand the interactions and structure of compact stars at such high density. However, the compact star EoS is constrained by several astrophysical and observational results for the structural properties like the mass, radius and tidal deformability. Such constraints include the data obtained from the most massive pulsar PSR J0740+6620 Fonseca et al. (2021); Miller et al. (2021); Riley et al. (2021), gravitational wave (GW170817) Abbott et al. (2018), PSR J0030+0451 (NICER)Riley et al. (2019); Miller et al. (2019), and HESS J1731-347 Doroshenko et al. (2022).
Many theoretical studies predict that strange quark stars (SQSs), composed of $u$, $d$ and $s$ quarks, may exist Olinto (1987). Such prediction is based on the Bodmer-Witten conjecture, which states that the energy per baryon density $\varepsilon/\rho_{B}$ of strange quark matter (SQM) is lower than that of pure nucleonic system Bodmer (1971); Chin and Kerman (1979); Witten (1984); Farhi and Jaffe (1984). Consequently, over many years, the theoretical modeling of SQM has been a topic of great interest. The first of such kind of models proposed is the well known MIT Bag model Chodos et al. (1974), which is later modified to ensure the repulsive effect between the quarks by introducing a vector meson as mediator (vBag model) Klahn and Fischer (2015); Cierniak et al. (2019); Gomes et al. (2019); Franzon et al. (2016); Wei et al. (2019); Lopes et al. (2021); Kumar et al. (2022). In the MIT Bag model, the difference in energy density between the perturbative vacuum and the true vacuum is represented by a constant $B$ whose value is still not well known. The stability of SQS in terms of $\varepsilon/\rho_{B}$ is found to be controlled by $B$ Farhi and Jaffe (1984); Torres and Menezes (2013) and further, the allowed range of $B$ for different vector meson couplings ($G_{V}$) is estimated in Ref. Lopes et al. (2021).
Several astrophysical and cosmological observations like the rotation curves of the galaxies, gravitational lensing, X-ray analysis of Bullet cluster Ade et al. (2016); Bertone et al. (2005); Aghanim et al. (2020); Bauer et al. (2020) support the fact that the dark matter (DM) to luminous baryonic matter ratio of the total energy budget in the Universe is around 5:1. The important experimental search avenues including the direct and the indirect detection strategies Lisanti (2017); Profumo (2013) are actively pursuing the intricacies of interaction between the DM and standard model particles. Of these searches, the most important direct detection experiments are the SuperCDMS Agnese et al. (2018), XENONnT Aprile et al. (2022), PandaX-II Wang et al. (2020), DarkSide-50 Agnes et al. (2018), SENSEI Crisler et al. (2018) and LUX-ZEPLIN Aalbers et al. (2023); while, the indirect searches are FERMI-LAT Atwood et al. (2009), IceCube Aartsen et al. (2014), PAMELA Adriani et al. (2009, 2013), AMS-02 Accardo et al. (2014), Voyager Boudaud et al. (2017) and CALET Adriani et al. (2017, 2018). However, the exact nature, properties and interaction of DM particle candidates are still largely unknown. The present day thermal relic abundances of DM has been prescribed to be $\sim\Omega_{DM}h^{2}\approx 0.12$ Tanabashi et al. (2018); Bauer and Plehn (2019); Cannoni (2016) from the Cosmic Microwave Background (CMB) anisotropy maps, obtained from the Wilkinson Microwave Anisotropy Probe (WMAP) data Bennett et al. (2013). Here, the relic abundances of DM ($\Omega_{DM}$) is conventionally estimated in terms of the critical density of the Universe $\rho_{crit}$ and the scale factor for the Hubble expansion rate $h$. Thus any reasonable DM model is constrained to successfully reproduce the observed non-baryonic relic density.
The compact stars like SQSs are extremely gravitating objects which are capable of accreting matter from their surroundings and such accreted matter may include DM Baryakhtar et al. (2017); Joglekar et al. (2020a); Bell et al. (2019); Joglekar et al. (2020b); Bramante and Raj (2024). Few other phenomena may also be attributed to the possible presence of DM inside compact stars; like, dark decays of neutrons Husain et al. (2022); Husain and Thomas (2023); Husain et al. (2023); Zhou (2023), inheritance of DM from the supernovae Nelson et al. (2019) etc. It is therefore reasonable to consider the DM particle density $\rho_{\chi}$ to be almost constant throughout the radial profile of the star Panotopoulos and Lopes (2017a); Guha and Sen (2021); Sen and Guha (2021, 2022); Guha and Sen (2024). DM thus becomes structural part of the SQSs, consequently allowing us to study the properties of DM admixed SQSs (DMSQSs) in the light of the different observational and astrophysical constraints on the structural properties of compact stars Yang and Pi (2024); Rezaei (2023); Panotopoulos and Lopes (2017b); Zheng and Chen (2016); Mukhopadhyay and Schaffner-Bielich (2016); Jiménez and Fraga (2022); Yang et al. (2021); Mafa Takisa et al. (2020); Panotopoulos and Lopes (2018a, b); Cassing et al. (2023); Ferreira and Fraga (2023). DM may or may not interact with SQM; however, if they interact, then the interaction strength must be extremely weak in order to prevent the star from collapsing Zheng and Chen (2016). Based on this notion, in Ref. Sen and Guha (2022), a feeble interaction between SQM and fermionic DM $\chi$ via a dark vector boson mediator $\xi$ was introduced. The pure SQM is described with the vBag model. The mass of DM fermion $m_{\chi}$ and the mass of vector mediator $m_{\xi}$ are chosen by following the self-interaction constraint $\sigma_{T}/m_{\chi}\leq 1.25~{}\rm{cm^{2}/g}$ Randall et al. (2008); Robertson et al. (2017) from the Bullet cluster Tulin et al. (2013); Tulin and Yu (2018); Hambye and Vanderheyden (2020). Here, $\sigma_{T}$ is the self-scattering transfer cross-section. Further, the coupling $y_{\xi}$ between $\chi$ and $\xi$ are obtained by satisfying the present day relic abundance of DM fermion $\chi$ Belanger et al. (2014); Gondolo and Gelmini (1991); Guha et al. (2019). In the present work, we follow the same treatment as in Ref. Sen and Guha (2022) for the interaction between DM and SQM, in order to investigate its effects on the non-radial oscillations of the DMSQSs.
Gravitational waves (GWs) can be produced by binary neutron star (BNS) mergers as well as from isolated neutron stars (NSs). Following the first detection of a GW signal from a binary black hole merger Abbott et al. (2016), many more signals have been detected from supernovae, NS mergers and supermassive black hole mergers. Furthermore, the NanoGrav collaboration recently revealed the existence of a cosmic gravitational wave background  Agazie et al. (2023); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023). The detection of GWs from a BNS merger (GW170817) Abbott et al. (2017) by the advanced LIGO and VIRGO collaborations combined with the multi-messenger observations of BNS systems have paved a new way for studying the interior composition of compact stars. The non-radial pulsations of compact stars are of current interest in the era of gravitational wave detection because they are related to the emission of GWs of different frequencies viz., fundamental ($f$), pressure ($p$), rotational ($r$), space-time ($w$) and gravity ($g$) modes Kokkotas and Schmidt (1999). There have been a lot of studies on these oscillation modes done over the past few years Lindblom and Owen (2002); Jha et al. (2010); Flores and Lugones (2014); Andersson and Pnigouras (2019); Jyothilakshmi et al. (2022); Zhao and Lattimer (2022); Kumar et al. (2024), among which $f$-modes have drawn a great interest since their frequencies ($1-3$ kHz) lie in the range of the future GW detectors such as the Einstein Telescope and the Cosmic Explorer and also the LIGO O4 run. With these next generation GW detectors the $f$-mode oscillations are expected to be observed; which will provide more powerful insights to the composition and interaction of matter inside compact stars and help in constraining the EoS in stringent manner.
For the purpose of gravitational wave asteroseismology, the estimation of the gravitational-radiation flux from compact stars is based on the theory of non-radial oscillations, which was initially developed in the framework of general relativity in Thorne and Campolattaro (1967); Lindblom and Detweiler (1983). Later, with the help of Cowling approximations Cowling (1941), the derivation of the integrated solution was simplified Sotani et al. (2011) by neglecting the metric perturbations. Although, this approximation can result in overestimation of the $f$-modes upto 30% Zhao and Lattimer (2022), it is widely used due to its ability to bring out qualitatively the essential characteristics of $f$-modes. Recently, there have been several studies on non-radial $f$-mode oscillations using the Cowling approximation for different stellar composition Sotani et al. (2011); Flores and Lugones (2014); Ranea-Sandoval et al. (2018); Vásquez Flores et al. (2019); Pradhan et al. (2022); Kalita et al. (2024); Zhang et al. (2024); Jyothilakshmi et al. (2024). Further, the non-radial $f$-mode oscillations of dark matter admixed NSs Das et al. (2021); Hong et al. (2023); Shirke et al. (2024); Sen and Guha (2024) were also studied. In the case of DMSQSs, although the radial oscillations have attracted the interest of the community recently Panotopoulos and Lopes (2017b, 2018a); Jiménez and Fraga (2022); Zhen et al. (2024); the non-radial oscillations have not been explored so far. This sets the motivation for our analysis. In the present work, we are particularly interested to show how the presence of DM and its interaction with SQM affects the non-radial $f$-mode frequency of the oscillation of the DMSQSs using the Cowling approximation.
We arrange the present work as follows. In the next Sec. II, we briefly discuss the mechanism of invoking feeble interaction of DM with SQM via the new physics vector mediator $\xi$. We also discuss the methodology of estimating the structural properties like mass, radius, tidal deformability and the $f$-mode oscillation of the DMSQSs. We then present the results and corresponding discussions in Sec. III. We summarize and conclude in the final section IV of the paper. Throughout this paper we set $c=G=1$, where $c$ is the speed of light in vacuum and $G$ is the universal gravitational constant.
II Formalism
II.1 Dark matter admixed quark star model
Here, we review the DMSQS model as derived in Ref. Sen and Guha (2022). The model consists of two main sectors: strange quark matter (SQM) and dark matter (DM), together with their interactions. The vector MIT Bag (vBag) model is employed to describe the $u$, $d$ and $s$ quarks along with the electrons $e$ Lopes et al. (2021); Kumar et al. (2022). The complete modified Lagrangian for such a system is given as Sen and Guha (2022)
$\mathcal{L}$ $=$ $\sum_{f}\bigg{[}\overline{\psi}_{f}\bigg{(}i\gamma_{\mu}\partial^{\mu}-g_{\xi}\gamma_{\mu}\xi^{\mu}-m_{f}\bigg{)}\psi_{f}-B\bigg{]}\Theta(\overline{\psi}_{f}\psi_{f})$ (1)
-- $\sum_{f}g_{qqV}\bigg{[}\overline{\psi}_{f}\bigg{(}\gamma_{\mu}V^{\mu}\bigg{)}\psi_{f}\bigg{]}\Theta(\overline{\psi}_{f}\psi_{f})-\frac{1}{4}V_{\mu\nu}V^{\mu\nu}+\frac{1}{2}m_{V}^{2}V_{\mu}V^{\mu}$
$+$ $\overline{\psi}_{l}\bigg{(}i\gamma_{\mu}\partial^{\mu}-m_{l}\bigg{)}\psi_{l}-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{1}{2}m_{\xi}^{2}\xi_{\mu}\xi^{\mu}$
$+$ $\overline{\chi}\bigg{[}i\gamma_{\mu}\partial^{\mu}-y_{\xi}\gamma_{\mu}\xi^{\mu}\bigg{]}\chi;$
where $f\equiv$ ($u$, $d$, $s$) and the lepton $l\equiv e$. The term $\overline{\psi}_{f}(g_{\xi}\gamma_{\mu}\xi^{\mu})\psi_{f}$ represents the interaction between the quark fields $\psi_{f}$ and the new physics vector dark mediator $\xi$ of mass $m_{\xi}$, with a very feeble coupling strength $g_{\xi}$. The Bag constant is denoted by $B$ and the Heaviside step function $\Theta$ in the Lagrangian ensures that the quarks are confined within the bag. The first term in second line denotes the interaction between quarks, mediated by the repulsive vector meson $V$ having mass $m_{V}$, with coupling strength $g_{qqV}$. Further, a universal coupling scheme is considered i.e., $g_{uuV}=g_{ddV}=g_{ssV}=g_{qqV}$ and $G_{V}=(g_{qqV}/m_{V})^{2}$ denotes the scaled coupling. Note that $G_{V}=0$ results in the original form of the MIT Bag model without interactions. The kinetic term of the vector meson in the SQM is given by $\frac{1}{4}V_{\mu\nu}V^{\mu\nu}$ with $V_{\mu\nu}=\partial_{\mu}V_{\nu}-\partial_{\nu}V_{\mu}$. Further, the kinetic term of the dark mediator $\xi$ is given as $\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$, where $F_{\mu\nu}=\partial_{\mu}\xi_{\nu}-\partial_{\nu}\xi_{\mu}$. The last line of Eq. (1) denotes the DM self interaction of fermionic $\chi$ involving bosonic dark mediator $\xi$, with coupling $y_{\xi}$. The interaction between the quarks $\psi_{f}$ and fermionic DM $\chi$, mediated by the vector dark boson $\xi$, can be understood with the help of the Feynman diagram (Fig. 1), which depicts the reversible process of $\chi\overline{\chi}\stackrel{{\scriptstyle\xi}}{{\rightleftharpoons}}{\psi}_{f}\overline{\psi}_{f}$.
$\psi_{f}$$\overline{\psi}_{f}$$\xi$$\chi$$\overline{\chi}$$g_{\xi}$$y_{\xi}$
$\chi$$\overline{\chi}$$\xi$$\psi_{f}$$\overline{\psi}_{f}$$y_{\xi}$$g_{\xi}$
Figure 1:  Feynman diagrams for the fermionic dark matter interaction with quarks via dark mediator, $\chi\overline{\chi}\stackrel{{\scriptstyle\xi}}{{\rightleftharpoons}}{\psi}_{f}\overline{\psi}_{f}$.
In this model, the stability window for Bag constant $B$ is fixed for a particular value of $G_{V}$, following the procedure described in Refs. Torres and Menezes (2013); Lopes et al. (2021). The upper bound of $B$ ($B_{max}$) is obtained from the Bodmer-Witten conjecture Bodmer (1971), which demands that the energy per baryon density $\varepsilon/\rho_{B}$ at the surface of the SQS is less than the binding energy of iron nucleus i.e., $\varepsilon/\rho_{B}\leq 930$ MeV Torres and Menezes (2013); Lopes et al. (2021); while the lower bound $B_{min}$ is obtained with two flavor QM. To obtain the EoS from Eq. (1), we obtain an average value for $B$ ($B_{av}$) between $B_{max}$ and $B_{min}$, for a fixed value of $G_{V}$. Here, we take $G_{V}=0.3,0.5$, and 0.7 fm2 and the corresponding values of $B$ have been shown in Table 1. Further, we take the masses of quarks to be $m_{s}=95$ MeV, $m_{u}=5$ MeV and $m_{d}$ = 7 MeV and mass of vector meson is $m_{V}=783$ MeV.
{tblr}
hlines, vlines, colspec = Q[c,m,wd=1.6cm] Q[c,m,wd=1cm] Q[c,m,wd=1cm] Q[c,m,wd=1cm], rowsep=1pt $G_{V}$ (fm2) & 0.3 0.5 0.7
$B_{min}^{1/4}$ (MeV) 138 134 131
$B_{min}^{1/4}$ (MeV) 148 143 139
Table 1: Stability window for Bag constant $B$ obtained for the vBag model, for different values of $G_{V}$ Sen and Guha (2022).
It is interesting to note that, one of the most important features of the particular DMSQS model Sen and Guha (2022) is that the DM parameters $m_{\chi}$, $m_{\xi}$, and $y_{\xi}$ are not independent of each other or arbitrary. The masses $m_{\chi}$ and $m_{\xi}$ are related by the self-interaction constraint from Bullet cluster Tulin et al. (2013); Tulin and Yu (2018); Hambye and Vanderheyden (2020). One can also find the graphical representation of the $m_{\chi}$ vs $m_{\xi}$ relation in Ref. Sen and Guha (2022). The relation between $m_{\chi}$ and $y_{\xi}$ is obtained from the present day relic abundance of DM Belanger et al. (2014); Gondolo and Gelmini (1991); Guha et al. (2019). Hence, a variation in $m_{\chi}$ implies corresponding variation of both $m_{\xi}$ and $y_{\xi}$. We tabulate the mass of fermionic DM $\chi$ and the corresponding values of mass of the dark mediator $m_{\xi}$ and interaction strength $y_{\xi}$ used in the present analysis in Table. 2. Note that the interaction strength between the dark mediator and quarks is considered to be feeble, $g_{\xi}\sim 10^{-4}$. Choosing the $g_{\xi}$ values less than $10^{-4}$ will not bring any further change to the EoS as well as the structural properties of DMSQS. Also, the values of $g_{\xi}>10^{-4}$ are unphysical because in that case the interaction cross-section will be in the ruled out region from direct detection experiments.
{tblr}
hlines, vlines, colspec = Q[c,m,wd=1.6cm] Q[c,m,wd=1cm] Q[c,m,wd=1cm] Q[c,m,wd=1cm] Q[c,m,wd=1cm] Q[c,m,wd=1cm] Q[c,m,wd=1cm], rowsep=2pt $m_{\chi}$ (GeV) & 25 50 75 100 150 200
$m_{\xi}$ (MeV) 50 60 20 10 6 4
$y_{\xi}$ 0.27 0.32 0.40 0.46 0.50 0.52
Table 2: Values of the mass of fermionic DM $m_{\chi}$ and the corresponding values of mass of dark mediator $m_{\xi}$ and interaction strength $y_{\xi}$ Sen and Guha (2022).
In the mean field approximation, the equation of motion of the vector field in terms of $V_{0}$ is given by
$V_{0}=\frac{g_{qqV}}{m_{V}^{2}}\rho;$ (2)
where, $\rho$ is the total quark density
$\rho=\langle\psi_{f}^{\dagger}\psi_{f}\rangle=(\rho_{u}+\rho_{d}+\rho_{s})=\frac{\gamma_{f}}{6\pi^{2}}\sum_{f}{{k_{F}^{3}}_{f}}.$ (3)
Here, ${k_{F}}_{f}$ denotes the Fermi momenta of the quarks and $\gamma_{f}=6$ is the degeneracy factor for quarks. The vacuum expectation value ($\xi_{0}$) of the vector dark boson is
$\xi_{0}=\frac{g_{\xi}\rho+y_{\xi}\rho_{\chi}}{m_{\xi}^{2}};$ (4)
where, the DM density is
$\rho_{\chi}=\frac{\gamma_{f}}{6\pi^{2}}{{{k_{F}^{3}}}_{\chi}}.$ (5)
The number density $\rho_{\chi}$ is assumed to be constant throughout the radial profile of the star. We note that this constant number density of fermionic DM is $\sim 1000$ times smaller than the average baryon number density of the SQM Panotopoulos and Lopes (2017a); Guha and Sen (2021); Sen and Guha (2021, 2022); Guha and Sen (2024). Thus the Fermi momentum of the DM fermions, ${k_{F}}_{\chi}$, is also constant over the radial profile of star. Unlike Ref. Sen and Guha (2022), in the present analysis, we consider three different values for ${k_{F}}_{\chi}=0.02,0.03,0.04$ GeV. The quark chemical potential in presence of DM is modified as
$\mu_{f}=\sqrt{k_{f}^{2}+m_{f}^{2}}+g_{qqV}V_{0}+g_{\xi}\xi_{0}.$ (6)
The calculated EoS for the DMSQSs is given as follows:
[…]
$P$ $=$ $\frac{1}{2}\frac{g_{qqV}^{2}}{m_{V}^{2}}\rho^{2}+\frac{1}{2}\frac{\Big{(}g_{\xi}\rho+y_{\xi}\rho_{\chi}\Big{)}^{2}}{m_{\xi}^{2}}+\frac{\gamma_{f}}{6\pi^{2}}\sum_{f}\int_{0}^{{k_{F}}_{f}}\frac{k_{f}^{4}~{}dk_{f}}{\sqrt{k_{f}^{2}+m_{f}^{2}}}$
$+$ $\frac{\gamma_{l}}{6\pi^{2}}\int_{0}^{{{k_{F}}}_{l}}\frac{{{k_{l}^{4}}}~{}dk_{l}}{\sqrt{k_{l}^{2}+m_{l}^{2}}}+\frac{\gamma_{l}}{6\pi^{2}}\int_{0}^{{{k_{F}}}_{\chi}}\frac{{{k_{\chi}^{4}}}~{}dk_{\chi}}{\sqrt{k_{\chi}^{2}+m_{\chi}^{2}}}-B,$
where $\varepsilon$ and $P$ are the energy density and pressure respectively. The presence of charge neutral DM do not affect the overall charge neutrality of the star and the EoS is computed by satisfying the charge neutrality and chemical potential equilibrium conditions of the star Glendenning (2000).
II.2 Structural properties and non-radial oscillations of dark matter admixed quark stars
Using the EoS of DMSQSs, we estimate the global properties like the gravitational mass ($M$) and the radius ($R$) of the DMSQSs. In this section, we describe the formalism used to study the non-radial oscillations of the DMSQSs. The metric for a static spherically symmetric star is given as
$ds^{2}=-e^{2\Phi(r)}dt^{2}+e^{2\lambda(r)}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2}.$ (9)
Here, $\Phi$ and $\lambda$ are the metric functions. The Einstein field equations are solved for the given metric for an ideal fluid to obtain the well-known Tolman-Oppenheimer-Volkoff (TOV) equations Tolman (1939); Oppenheimer and Volkoff (1939) and is given as
$\frac{dP(r)}{dr}$ $=$ $-\Big{(}\varepsilon(r)+P(r)\Big{)}\frac{d\Phi(r)}{dr},$ (10)
$\frac{d\Phi(r)}{dr}$ $=$ $\frac{M(r)+4\pi r^{3}P(r)}{r\Big{(}r-2M(r)\Big{)}},$ (11)
$\frac{dM(r)}{dr}$ $=$ $4\pi r^{2}\varepsilon(r).$ (12)
These coupled differential equations are integrated from the center ($r=0$) to the surface ($r=R$) of the star. Towards the center of the star the pressure and mass of the star are $P_{c}=P(r=0)=P(\varepsilon_{c})$ and $M_{c}=M(r=0)=0$, respectively. The pressure approaches zero towards the surface of the star. By solving the TOV equations for all possible values of $\varepsilon_{c}$, we obtain the mass $M$ and radius $R$ of the star. The mass function $M(r)=r(1-e^{-2\lambda(r)})/2$ satisfies Eq. (12). Compactness ($C$) is a global parameter that quantifies the stellar gravity and it is defined in terms of the stellar mass $M$ and radius $R$ as $C=M/R$.
We adopt the methodology to calculate the non-radial oscillations of the DMSQSs using the Cowling approximations, which is well presented in Ref. Sotani et al. (2011). After solving the TOV Eqs. (10) - (12), we obtain the oscillations mode frequencies by solving two following coupled differential equations:
$\frac{dW(r)}{dr}$ $=$ $\frac{d\varepsilon(r)}{dP(r)}\Bigg{[}\omega^{2}r^{2}e^{\lambda(r)-2\Phi(r)}V(r)+\frac{d\Phi(r)}{dr}W(r)\Bigg{]}-l(l+1)e^{\lambda(r)}V(r),$ (13)
$\frac{dV(r)}{dr}$ $=$ $2\frac{d\Phi(r)}{dr}V(r)-e^{\lambda(r)}\frac{W(r)}{r^{2}}.$ (14)
In order to solve the Eqs. (13) and (14), the two boundary conditions at the center ($r=0$) and the surface ($r=R$) of the star are to be imposed. Near the center ($r=0$) of the star the functions $W(r)$ and $V(r)$ behave as
$W(r)=Ar^{l+1}~{}{\rm{and}}~{}~{}V(r)=-Ar^{l}/l;$ (15)
where $A$ is an arbitrary constant. Towards the surface ($r=R$) of the star the vanishing perturbation functions provide another boundary condition for $W(r)$ and $V(r)$ and is written as
$\omega^{2}R^{2}e^{\lambda(R)-2\Phi(R)}V(R)+\frac{d\Phi(r)}{dr}\Bigg{\rvert}_{r=R}W(R)=0.$ (16)
The coupled differential equations, Eqs. (13) and (14), are integrated from the center to the surface of the star by assuming an initial value of $\omega^{2}$. After each integration the value of $\omega^{2}$ is improved using Ridder’s method until Eq. (16) is satisfied.
The tidal deformability is an astrophysical observable that defines the quadrupole deformation of a compact star in a binary system due to the tidal field created by its companion star. In case of an isolated star, the tidal deformability is calculated in the limit where source of the static external quadrupolar tidal field is very far away Hinderer (2008). The dimensionless tidal deformability ($\Lambda$) is given as
$\Lambda=\frac{2}{3}k_{2}R^{5}.$ (17)
Here $k_{2}$ is the tidal love number, which is given in terms of the compactness ($C$), and a quantity ($y$) obtained by following the formalism developed in Ref. Hinderer (2008); Hinderer et al. (2010). In case of QSs, $y$ is defined as Hinderer et al. (2010); Kumar et al. (2022)
$y=\frac{RH^{\prime}(R)}{H(R)}-\frac{4\pi R^{3}\varepsilon_{s}}{M(R)};$ (18)
where, $\varepsilon_{s}$ is the energy density at the surface of the QS.
III Results
In this section, we discuss the results obtained for $f$-mode oscillations using the Cowling approximation (given in Sec. II.2) for DMSQSs (given in Sec. II.1). First, we compute the DMSQSs EoS using Eqs. (II.1) and (LABEL:P). With the DMSQSs EoS, we then obtain the static stellar structure by numerically solving the TOV equations given by Eqs. (10) to (12) from the center to the surface of the star. We then numerically integrate the coupled differential Eqs. (13) and (14) from the center to the surface of the star using the appropriate boundary conditions given by Eqs. (15) and (16). The $f$-mode frequencies are obtained with the required precision using Ridder’s method. We also calculate the tidal deformability $\Lambda$ of DM admixed QSs using the Eqs. (18) and (17). In order to study the effect of DM for different values of the SQM vector coupling constant $G_{V}$, we vary the DM particle mass $m_{\chi}$ and the Fermi momentum of the DM fermions ${k_{F}}_{\chi}$ and study their effects on the stellar structure and $f$-mode oscillation of the DMSQSs. We have investigated the explicit dependence of the DMSQS properties on $m_{\chi}$ and ${k_{F}}_{\chi}$ by varying them individually. While varying $m_{\chi}$ and ${k_{F}}_{\chi}$ we have kept the vBag parameters like $G_{V}$ and corresponding value of $B_{av}$ fixed. For example, we fixed $G_{V}$=0.3 fm2 and the corresponding value of $B_{av}$ when we varied $m_{\chi}$. Again we fix $G_{V}$=0.5 fm2 with the corresponding value of $B_{av}$ while varying ${k_{F}}_{\chi}$. While varying $m_{\chi}$, we have fixed ${k_{F}}_{\chi}$ and vice versa. This way, the dependence of the properties of DMSQS on both the mass and the number density of DM are examined explicitly and individually in the present paper. While varying the parameters of dark sector, the parameters of the MIT vbag model are kept fixed in order to study the distinct role that DM plays on the properties of DMSQS.
(a)  Fixed ${k_{F}}_{\chi}$= 0.03 GeV and different $G_{V}$ and $m_{\chi}$.
(b)  Fixed $G_{V}$ = 0.5 fm2 and $m_{\chi}$= 75 GeV and different ${k_{F}}_{\chi}$.
Figure 2: Equation of state of dark matter admixed strange quark star.
(a)  Mass profiles by varying $G_{V}$, ${k_{F}}_{\chi}$, and $m_{\chi}$.
(b)  Maximum mass region for different parameterizations by varying $G_{V}$, ${k_{F}}_{\chi}$, and $m_{\chi}$. We follow Fig. 2 for labels.
Figure 3: Variation of mass with radius of dark matter admixed strange quark stars (DMSQSs). The observational limits imposed on maximum mass from the most massive pulsar PSR J0740+6620 Fonseca et al. (2021) and corresponding radius Miller et al. (2021); Riley et al. (2021) are also indicated. The constraints on $M-R$ plane prescribed from GW170817 Abbott et al. (2018), the NICER experiment for PSR J0030+0451 Riley et al. (2019); Miller et al. (2019), and HESS J1731-347 Doroshenko et al. (2022) are also compared. The mass of the secondary component of GW190814 Abbott et al. (2020) is also displayed.
(a)  Fixed ${k_{F}}_{\chi}$= 0.03 GeV and $G_{V}$= 0.3 fm2 and different $m_{\chi}$.
(b)  Fixed $G_{V}$= 0.5 fm2 and $m_{\chi}$= 75 GeV and different ${k_{F}}_{\chi}.$
Figure 4: Variation of compactness $C=M/R$ with radius of DMSQSs.
We first study the effects of variation of $m_{\chi}$ on the EoS and structural properties of the DMSQSs for a fixed value of ${k_{F}}_{\chi}$=0.03 GeV and by considering different values of $G_{V}$. Later, we will investigate the effects of variation of ${k_{F}}_{\chi}$ for fixed $G_{V}$ and $m_{\chi}$ values. The EoS of the DMSQSs for different values of $m_{\chi}$= 50 and 100 GeV and $G_{V}$= 0.3 and 0.7 fm2 are shown in Fig. 2(a). We can see that the higher values of $G_{V}$ lead to stiffer EoS and consequently massive SQSs with larger radius (see also Fig. 3). This trend is similar to that seen in Ref. Lopes et al. (2021). So, for any particular value of $m_{\chi}$, stronger coupling strength between quarks gives massive DMSQSs with larger radius. Also, for any fixed value of $G_{V}$, Fig. 2(a) shows that the EoS softens with higher values of $m_{\chi}$. This is also reflected in the structural properties of the DMSQSs as seen in Fig. 3. We can see that the mass and radius of the DMSQSs reduce for massive DM. The reason can be attributed to the fact that presence of DM makes the star more compact and the heavier fermionic DM makes the star even more compact. To illustrate this, we show the variation of dimensionless compactness ($C$) with respect to radius in Fig. 4(a) for $G_{V}$=0.3 fm2. It can be inferred that, for a particular value of radius, $C$ increases with $m_{\chi}$ in order to incorporate massive DM at constant ${k_{F}}_{\chi}$ or constant $\rho_{\chi}$. Thus, lighter DM yields massive but less compact DMSQSs. For a low value of $m_{\chi}$, not only the mass but also the radius of the DMSQS increase which makes the star less compact. For example, the combination $G_{V}$=0.3 fm2 and $m_{\chi}$=50 GeV can produce a 1.4 $M_{\odot}$ DMSQS of corresponding radius $R_{1.4}$=11.60 km and corresponding compactness $C$=0.178; while $m_{\chi}$=100 GeV produces a 1.4 $M_{\odot}$ DMSQS of radius $R$=11.30 km and corresponding compactness $C$=0.183.
Next, in Fig. 2(b), we present the EoS of DMSQSs for different values of ${k_{F}}_{\chi}$ for fixed $G_{V}$ = 0.5 fm2 and $m_{\chi}$= 75 GeV. We notice that, with $G_{V}$=0.5 fm2, $m_{\chi}$=75 GeV, the EoS softens with increasing values of ${k_{F}}_{\chi}$ or $\rho_{\chi}$. This is because if the QS is largely populated by fermionic DM, the compactness $C$ increases as seen from Fig. 4(b), leading to softening of the EoS and consequently decreasing the mass and radius of the star, as seen from Fig. 3. For example, with fixed $G_{V}$=0.5 fm2 and $m_{\chi}$=75 GeV we obtain a 1.4 $M_{\odot}$ DMSQS of radius $R_{1.4}$=11.12 km and corresponding compactness $C$=0.186 for ${k_{F}}_{\chi}$=0.04 GeV but ${k_{F}}_{\chi}$=0.02 GeV produces a 1.4 $M_{\odot}$ DMSQS of radius $R$=12.24 km and corresponding compactness $C$=0.169. So, SQSs that are largely populated with massive DM fermions, are highly compact with comparatively less masses and radii.
We also find from Fig. 3(a) that the various astrophysical constraints on the $M-R$ relation of compact stars from the most massive PSR J0740+6620, GW170817, the NICER data for PSR J0030+0451, and HESS J1731-347 are well satisfied with the DMSQS EoS. We also observe that the mass of the secondary component of GW190814 Abbott et al. (2020) is also satisfied by the SQSs and with $G_{V}>$ 0.3 fm2 for the DMSQS. However, the exact nature of the massive secondary component of GW190814 is not known; whether it is a NS/QS or a black hole. As also seen in Ref. Sen and Guha (2022), all these constraints are better satisfied for higher values of $G_{V}$, since in Fig. 3(a), we find that for $G_{V}$=0.3 fm2 the NICER data for PSR J0030+0451 are violated for $m_{\chi}$=100 GeV. However, in Ref. Sen and Guha (2022), the recently obtained constraint from HESS J1731-347 Doroshenko et al. (2022) was not included. Therefore, in the present work we also test the DMSQSs EoS in the light of this constraint obtained from the low mass pulsar with low radius. We find that this constraint is successfully satisfied with our compact DMSQSs configurations. On the other hand, for $G_{V}$=0.5 fm2, $m_{\chi}$=75 GeV, shows that low DM content with low ${k_{F}}_{\chi}$ or $\rho_{\chi}$ can better satisfy the observational constraints since ${k_{F}}_{\chi}$= 0.04 GeV makes the star too compact to satisfy the NICER data for PSR J0030+0451. We also show the maximum mass region for various parameterisations of DMSQS in Fig. 3(b). Here, we have included all the configurations discussed in this work and it follows the label given in Fig. 2. As discussed earlier, an increase in the value of $G_{V}$ results in a stiffer EoS and this results in a larger maximum mass value. Similarily, we also find that with an increase in ${k_{F}}_{\chi}$ and $m_{\chi}$ values, the EoS becomes less stiff and results in stellar profiles with lower maximum mass values.
(a)  Fixed ${k_{F}}_{\chi}$= 0.03 GeV and different $G_{V}$ and $m_{\chi}$.
(b)  Fixed $G_{V}$= 0.5 fm2 and $m_{\chi}$= 75 GeV and different ${k_{F}}_{\chi}$.
Figure 5: Variation of $f$-mode frequency with mass of dark matter admixed strange quark star (DMSQS).
(a)  Fixed ${k_{F}}_{\chi}$= 0.03 GeV and different $G_{V}$ and $m_{\chi}$.
(b)  Fixed $G_{V}$= 0.5 fm2 and $m_{\chi}$= 75 GeV and different ${k_{F}}_{\chi}$.
Figure 6: Variation of $f$-mode frequency with compactness of DMSQS.
We now extend our work to focus on the $l=2$ $f$-mode oscillations of DMSQSs in Fig. 5, where we plot the $f$-mode frequencies as a function of stellar mass by fixing the value of the coupling constant $G_{V}$=0.3, 0.7 fm2 and the Fermi momentum of DM fermions ${k_{F}}_{\chi}=0.03$ GeV (Fig. 5(a)) for $m_{\chi}$= 50 and 100 GeV. We find that, the $f$-mode frequencies of DMSQSs decrease slightly with the mass and it only starts to increase towards the maximum mass of the star. With increase in value of $G_{V}$, we observe an overall decrement in the $f$-mode spectra of DMSQS. For any value of $G_{V}$, the $f$-mode frequencies show an overall increment with increase in the value of $m_{\chi}$. In the ‘no-DM’ (pure SQS) scenario, we find that larger vector repulsion between the quarks leads to lowering of the $f$-mode frequency of QSs. For example, a 1.4 $M_{\odot}$ SQS with $G_{V}$= 0.3 fm2 has $f$-mode oscillation frequency as high as 2.05 kHz compared to that with $G_{V}$= 0.7 fm2 having $f$= 1.81 kHz. In the presence of DM, lighter fermionic DM yields more massive DMSQSs with lower value of $f$-mode oscillation frequency. For example, with $G_{V}$=0.7 fm2, the 1.4 $M_{\odot}$ DMSQS produced for $m_{\chi}$=50 GeV oscillates with frequency $f$=1.92 kHz but $m_{\chi}$=100 GeV yields a 1.4 $M_{\odot}$ DMSQS with oscillation frequency $f$=2.02 kHz. Next, in Fig. 5(b), we plot the $f$-mode oscillations of DMSQSs by fixing $G_{V}=0.5$ fm2 and $m_{\chi}=75$ GeV, for different values of ${k_{F}}_{\chi}=0.02,0.03$ and 0.04 GeV. We observe that the $f$-mode spectra are more sensitive to the variations in ${k_{F}}_{\chi}$, compared to the changes in $m_{\chi}$ values. It can be seen that, larger values of ${k_{F}}_{\chi}$ results in overall increment in $f$-mode oscillations, compared to the no-DM case. For a 1.4 $M_{\odot}$ DMSQS, with ${k_{F}}_{\chi}$= 0.02 (0.04) GeV, we have $f_{1.4}$= 1.99 (2.09) kHz.
We now study the explicit dependence of $f$-mode frequency of DMSQSs on the compactness of the star in Fig. 6. As we have already seen that the presence of DM largely affects the compactness of the SQSs (Fig. 4), it is expected that this will be reflected in $f$-modes. For any fixed value of $C$, the $f$-mode frequency is found to be higher for lower coupling constant $G_{V}$ and for higher mass values of fermionic DM $\chi$ (see Fig. 6(a)). Similar to $m_{\chi}$, an increment in ${k_{F}}_{\chi}$ results in a more sensitive enhancement of $f$-modes as shown in Fig. 6(b). Here, we conclude that, highly compact DMSQSs, having massive DM (Fig. 6(a)) or high DM content in terms of high ${k_{F}}_{\chi}$ (Fig. 6(b)), oscillate faster with higher values of $f$-mode oscillation frequency.
(a)  Tidal deformability values of DMSQS by varying $G_{V}$, ${k_{F}}_{\chi}$, and $m_{\chi}$.
(b)  Tidal deformability values of DMSQS near to $1.4M_{\odot}$ profile. For labels refer Fig. 2.
Figure 7: Variation of tidal deformability with mass of DMSQS. The constraints on $\Lambda_{1.4}$ from GW170817 Abbott et al. (2018) and GW190814 Abbott et al. (2020) are also shown.
(a)  Fixed ${k_{F}}_{\chi}$= 0.03 GeV and different $G_{V}$ and $m_{\chi}$.
(b)  Fixed $G_{V}$= 0.5 fm2 and $m_{\chi}$= 75 GeV and different ${k_{F}}_{\chi}$.
Figure 8: Variation of $f$-mode frequency with tidal deformability of dark matter admixed strange quark star (DMSQS).
We plot the tidal deformability $\Lambda$ as a function of stellar mass by varying $G_{V}$, $m_{\chi}$, and ${k_{F}}_{\chi}$ values in Fig. 7(a). We find that the value of $\Lambda$ decreases with an increase in stellar mass till it attains the maximum mass. Also, in Fig. 7(b), we show the tidal deformability values close to $1.4M_{\odot}$ ($\Lambda_{1.4}$) star along with the constraints given by GW170817 and GW190814. We have included all the parameterizations of DMSQSs discussed in this work and it follows the labels given in Fig. 2. We find that the constraint on $\Lambda_{1.4}$ from GW170817 is satisfied with $m_{\chi}$=100 GeV and $G_{V}$=0.3 fm2. For the no-DM case as well as the DMSQSs with $m_{\chi}$=50 GeV, this constraint is not at all satisfied for any values of $G_{V}$ because we obtain very massive DMSQS configurations whose tidal deformability is also very high, thereby violating this constraint. However, the constraint from GW190814 on $\Lambda_{1.4}$ is comparatively better satisfied for the parameters considered for DMSQSs in this work. In Fig. 7(a), we find that this constraint is well satisfied by both SQS and DMSQSs for $G_{V}$=0.3 fm2. However, if there is strong repulsion among the quarks via high vector quark coupling strength then we require the presence of DM, rather massive DM to satisfy this constraint. The combined effect of the strong interaction between the quarks balanced by the feeble SQM-DM interaction helps to reduce the mass, radius and hence the tidal deformability of the DMSQSs, thereby helping to satisfy the constraint from GW190814 on $\Lambda_{1.4}$. We also study the relation between tidal deformability and mass of the SQSs for the variation of ${k_{F}}_{\chi}$ for fixed $G_{V}$ and $m_{\chi}$. We notice that in order to satisfy the constraints on $\Lambda_{1.4}$ from GW190814, high content of DM is necessary by the SQSs. Moderately massive ($m_{\chi}$ = 75 GeV) DM of Fermi momentum as high as ${k_{F}}_{\chi}$= 0.04 GeV is also capable of satisfying the constraint from GW170817 on $\Lambda_{1.4}$. However, we have seen from Fig. 3(a) that this combination of ($G_{V}$, $m_{\chi}$, ${k_{F}}_{\chi}$)=(0.5 fm2, 75 GeV, 0.04 GeV) do not satisfy the NICER data for PSR J0030+0451. So an intermediate value of ${k_{F}}_{\chi}$ between 0.03 and 0.04 GeV may satisfy both the constraints from the NICER data for PSR J0030+0451 as well as that on $\Lambda_{1.4}$ from GW170817.
(a)   Linear relation for $f$-mode oscillations and average density.
(b)   Angular frequency ($\omega=2\pi f$) normalized by stellar mass ($M$) as a function of compactness $C$.
Figure 9: (a) Variation of $f$-mode frequency with average density $\bar{\rho}=\sqrt{\bar{M}/\bar{R}^{3}}$ of DMSQSs. The solid red line shows the linear fit obtained for different parameterizations of DMSQSs used in this work. We also plot the linear fit obtained by Das et al. Das et al. (2021) denoted by (DKBP) and that obtained in Benhar et al.  Benhar et al. (2004) represented with (BFG), for comparison. (b) Variation of angular frequency $\omega$ normalised by stellar mass $M$ with compactness $C$. The solid red line represents the universal relation obtained for DMSQSs.
Further, in Fig. 8, we analyse the variation of the $f$-mode frequency of the DMSQSs with respect to its tidal deformability ($\Lambda$). We observe that, the mode frequencies slightly decrease with decrement in $\Lambda$ and it only begins to increase towards the $\Lambda$ corresponding to maximum mass of the star. We plot the $f$-modes for fixed ${k_{F}}_{\chi}$ and different $G_{V}$, $m_{\chi}$ in Fig. 8(a). As shown in Figs. 5(a) and 6(a), the $f$-mode frequencies decrease for high values of $G_{V}$; while for a fixed $G_{V}$, the frequencies increase with increment in $m_{\chi}$. In Fig. 8(b), we plot the $f$-modes as a function of $\Lambda$ for fixed $G_{V}$, $m_{\chi}$ and by varying ${k_{F}}_{\chi}$. As observed in Figs. 5(b) and 6(b), $f$-mode oscillations are sensitive to variation in ${k_{F}}_{\chi}$ and large values of ${k_{F}}_{\chi}$ results in higher values of frequencies.
Finally, for different parameterizations of DMSQSs, we obtain a linear (empirical) relation between the $f$-mode frequency and the average density $\bar{\rho}=\sqrt{\bar{M}/\bar{R}^{3}}$, where $\bar{M}=M/(1.4$ $M_{\odot}$) and $\bar{R}=R/(10$ km). The empirical formula of the form, $f=a+b\bar{\rho}$ was first obtained by Andersson and Kokkotas Andersson and Kokkotas (1996) and by knowing the values of $f$-mode frequencies, we can calculate the mass and radius of the star. This could therefore help in constraining the EoS more closely Andersson and Kokkotas (1996). For our EoSs, we obtain the values of $a$ and $b$ as 1.367 kHz and 0.87 kHz respectively. In Fig. 9(a), we plot the relation obtained for DMSQSs along with the empirical relations from relevant earlier works for stellar models with dark matter admixed hyperon stars (DKBP) Das et al. (2021) and hybrid stars (BFG) Benhar et al. (2004). We find that our fit relation differ considerably from the already existing relations from DKBP and BFG. However, our fit is found to be more close to the fit for hybrid stars denoted by BFG. Further, we also obtain the correlation between the angular frequency scaled by mass ($\omega M$) and the compactness ($C$) of the DMSQSs. In Ref. Wen et al. (2019), Wen et al. found that the variation of $\omega M$ as a function of $C$ is universal for $f$-modes. Similarly, in Fig. 9(b), we fit the mass scaled $\omega$ with $C$ for different parameterizations of DMSQSs. The universal relation was found to take the form $\omega M$ (kHz km) $=129.50\,(M/R)-3.48$, where the coefficients are obtained after performing a linear fit. In order to obtain these correlations we varied the values of $G_{V}$ from 0.3 to 0.7 GeV and $m_{\chi}$ from 50 to 100 GeV.
IV Summary and Conclusion
In the present work, we have investigated the non-radial $l=2$ $f$-mode oscillations of strange quark stars (SQSs) admixed with dark matter (DM). For our analysis, we have considered the presence of DM in SQSs, with the interaction between the fermionic DM $\chi$ and SQM $\psi_{f}$ being mediated by the dark mediator $\xi$. We note that, to the best of our knowledge, this is the first time the non-radial $f$-mode spectra of DMSQSs have been studied.
Firstly, we looked into the structural properties of DMSQSs. The EoS of DMSQSs softens with both massive and higher population of fermionic DM and as a result, we obtain massive DMSQS configurations with larger radius for lower values of $m_{\chi}$ and ${k_{F}}_{\chi}$. This is because the presence of DM makes the star more compact and this compactness is further individually enhanced by heavier and dense population of fermionic DM. Overall, very compact and massive DMSQS configurations, that not only satisfy the constraints from PSR J0740+6620 on the mass-radius relation of compact stars but also the mass of the secondary component of GW190814, are achieved. Many of the DMSQS configurations also satisfy the various recent observational and astrophysical constraints from GW170817, the NICER experiment for PSR J0030+0451, and HESS J1731-347.
The prominent non-radial $f$-mode frequencies of the DMSQSs are calculated by considering the general relativistic Cowling approximation. By obtaining the $f$-modes as a function of stellar mass, compactness and tidal deformability, we have studied in detail how the presence of DM and its interaction with SQM affect the $f$-mode spectra. We have found that the $f$-mode frequencies of DMSQSs, which are largely populated with massive DM fermions, are higher compared to that of the no-DM scenario (SQSs). We also obtained that the fundamental $f$-mode oscillations are more sensitive to the variations in Fermi momentum of DM ${k_{F}}_{\chi}$, than to the changes in its mass $m_{\chi}$. Thus the presence and interaction of DM with SQM has profound impact on the structural and oscillation properties of the DMSQSs. We conclude that the presence of massive DM of higher density resulted in comparatively less massive, but very compact DMSQSs with larger $f$-mode oscillation frequencies. An empirical relation between the $f$-modes and the average density $\bar{\rho}=\sqrt{\bar{M}/\bar{R}^{3}}$ was also determined and we found that the linear fit for DMSQSs differ significantly from the earlier works. We also estimated relations between the mass-scaled angular frequency $\omega M$ with the compactness $C$. It was observed that $\omega M$ varies universally with $C$ for DMSQSs.
Acknowledgements
L.J.N. acknowledges the Department of Science and Technology, Government of India for the INSPIRE Fellowship. Work of DS is supported by the NRF research Grants (No. 2018R1A5A1025563). Work of AG is supported by the National Research Foundation of Korea, grant funded by Korea Government (MSIT) (RS-2024-00356960).